OpenStudy (anonymous):

Help pretty please:) Will give medal. Slove:d+3/8d-2d+1/10d^2 A.)12d-4/40d^2 B.)12d^2-4/40d^2 C.)5d^2+7d-4/40d^2

OpenStudy (anonymous):

\[\huge \frac{d+3}{\frac{8d-2d+1}{10d^2}}\]?

OpenStudy (anonymous):

is there a typo there maybe? this seems unlikely

OpenStudy (anonymous):

ooh maybe it is \[\huge \frac{d+3}{8d}-\frac{2d+1}{10d^2}\] i get that is it

OpenStudy (anonymous):

common denominator is \(40d^2\)

OpenStudy (anonymous):

numerator is \[4(d+3)-5d(2d+1)=-(10 d^2+d-12)\]

OpenStudy (anonymous):

thank you